OpenStudy (anonymous):

how to solve (h*j)=h(x) * j(x) h(x)=4x^2 j(x)=6x^3 the multiplication symbol (*) is a closed circle

OpenStudy (amistre64):

what is 4x^2 times 6x^3 ???

OpenStudy (amistre64):

open circle is a nesting symbol; closed circle can either mean a dot product (vectors) or just multiplication

OpenStudy (anonymous):

the carrot symbol (^) means it's raising it to the whatever power

OpenStudy (amistre64):

yes, ^ is rather standard for exponent

OpenStudy (anonymous):

so x^3 means x is raised to the third power

OpenStudy (amistre64):

\[(4xx)(6xxx)=4(6)(xxxxx)=24x^5\]

OpenStudy (anonymous):

If the circle is ° but centered, it may be (he function composition operator (eg, f°g(x)= f(g(x)) . In this case we may have to find x such as h°j(x) = h(x) * j(x). One trivial solution is 0.

OpenStudy (anonymous):

okay so you add the exponents and don't multiply them..right?

OpenStudy (amistre64):

adding exponents is correct. when in doubt, just spread them out x^2 = x*x x^3 = x*x*x x^2 * x^3 = x*x*x*x*x = x^5

OpenStudy (anonymous):

24x^5

OpenStudy (anonymous):

so for example, for (2x-1)(4x^2), would the answer be 8x^3-4x^2?

OpenStudy (amistre64):

yes